# Generation and descent for derived categories of noncommutative structure sheaves

### v2

@@ Abstract.
This work demonstrates classical generation is preserved by the derived pushforward along the canonical morphism of a noncommutative structure sheaf to its underlying scheme. There are intriguing examples illustrating this phenomenon, particularly from noncommutative resolutions, categorical resolutions, and homological projective duality. Additionally, we establish that the Krull dimension of a variety over a field is a lower bound for the Rouquier dimension of the bounded derived category associated with a noncommutative structure sheaf on it. This is an extension of a classical result of Rouquier to the noncommutative context.
@@
Derived categories, strong generation, Rouquier dimension, noncommutative structure sheaves, noncommutative crepant resolutions, categorical resolution of singularities, homological projective duality, Azumaya algebra
@@ 1. Introduction
In this note, we examine the behavior of generation along the derived pushforward of the canonical morphism from a noncommutative structure sheaf to its underlying scheme. As a result, our investigation introduces new methods coming from noncommutative homological techniques for establishing upper bounds on the Rouquier dimension within triangulated categories relevant to the commutative world.
@@
There has been recent advancements that illuminate sufficient conditions under which $D^{b}_{\operatorname{coh}}(X)$ possesses such objects: for strong generators, any Noetherian quasi-excellent separated scheme of finite Krull dimension [Aok21]; for classical generators, any Noetherian $J\textrm{-}2$ scheme [ELS20]. Albeit being aware of the existence of such objects, explicit descriptions are limited to a few instances: smooth quasi-projective schemes over a field [Rou08], Frobenius pushforwards on a compact generator for Noetherian schemes of prime characteristic [${}^{+}$23], singular varieties that admit a resolution of singularities [Lan23, DL24], and objects arising from module categories [DLT23].
@@
Recent attention has been drawn to the case where $X$ is a projective curve [BD11, BDG17b, HH23, BD17] with significant progress made in establishing upper bounds on the Rouquier dimension of its bounded derived category. In particular, [BDG17b] identified an essentially surjective functor $D^{b}_{\operatorname{coh}}(\mathbb{X})\to D^{b}_{\operatorname{coh}}(X)$, where $\mathbb{X}$ is a noncommutative projective curve, yielding a categorical resolution of $X$. In the context of this work, a noncommutative structure sheaf is defined as a pair $(X,\mathcal{A})$, where $X$ is a Noetherian scheme, and $\mathcal{A}$ is a coherent $\mathcal{A}$-algebra [YZ06, BDG17a]. Additional background information can be found in Section 3.1. For alternative approaches to noncommutative algebraic geometry, refer to [Orl16, KR00, Lau03, ${}^{+}$16].
In a closely related context, [ELS20] demonstrated that the bounded derived category of coherent $\mathcal{A}$-modules, denoted $D^{b}_{\operatorname{coh}}(\mathcal{A})$, admits a classical generator whenever $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra over a Noetherian $J\textrm{-}2$ scheme. Our work is motivated by two key aspects: the preservation of generation along a derived pushforward of a proper surjective morphism and the geometry of a noncommutative structure sheaf. This represents a blend of ideas from [Lan23, DL24, ELS20, BDG17a] and leads us to our first result: an investigation into the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative structure sheaf to its underlying scheme. The following is Theorem 3.1.
@@
There are interesting examples where Theorem A can be studied, as explored in more detail in Section 4. In particular, these include noncommutative resolutions (NCR), categorical resolution of singularities, and homological projective duality. These examples not only showcase interesting applications of noncommutative structure sheaves but also provide valuable strategies offering upper bounds for Rouquier dimension within the commutative setting. Furthermore, we introduce a construction that yields examples of NCRs, where the obtained bounds indicate that the Rouquier dimension is no larger than the global dimension of a specific endomorphism ring. This is illustrated in Lemma 4.2.3 and discussed further in Remark 4.2.4. The subsequent Corollary 3.2 encapsulates the results presented below.
@@ 2. Generation
    [${}^{+}$23, Corollary 3.9] If $X$ is a Noetherian $F$-finite scheme, $G$ is a compact generator, and $e\gg 0$, then $F_{\ast}^{e}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
@@ 3. Derived descent
In this section, we investigate the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative structure sheaf to its underlying scheme. The subsequent is our main result.
@@
Before delving into the proofs of our results, we take a brief detour to explore the essential background and machinery related to noncommutative structure sheaves.
3.1. Noncommutative structure sheaves
@@
    A noncommutative structure sheaf is a pair $(Y,\mathcal{B})$ where $Y$ is a scheme and $\mathcal{B}$ is a quasi-coherent sheaf of $\mathcal{O}_{Y}$-algebras.
@@
If $X$ is a scheme, then the pair $(X,\mathcal{O}_{X})$ is a noncommutative structure sheaf. Assume $\mathcal{A}$ is a quasi-coherent $\mathcal{O}_{X}$-algebra. Consider the morphism of noncommutative structure sheaves
@@
Let $X$ be a variety over a field. Consider a sheaf of Azumaya algebras $\mathcal{A}$ on $X$, i.e., a locally finite free sheaf on $X$ that is étale locally a matrix algebra. This sheaf corresponds to an element in the Brauer group of $X$. For a treatment of the functorial properties of the noncommutative structure sheaf $(X,\mathcal{A})$ and its bounded derived category, refer to [Kuz06, Appendix D]. The pair $(X,\mathcal{A})$ is called a Azumaya variety.
@@ 3.2. Results
To exhibit Theorem 3.1, we start by checking for exactness of the canonical morphism of a noncommutative structure sheaf to its underlying scheme. The following statement might be clear to experts, but we include it for the sake of coherence.
@@
[…]
@@ 4.4. Homological projective duality
Suppose $V$ is a $k$-vector space of odd dimension $v$. For each even number $0\leq 2q<v$, we have a Pfaffian variety $\operatorname{Pf}_{q}$ contained in $\mathbb{P}(\wedge^{2}V^{\vee})$ consisting of all $2$-forms on $V$ whose rank is at most $2q$. This variety is not a complete intersection and usually is singular. The projective dual of $\operatorname{Pf}_{q}$ is another Pfaffian variety whose locus is the subspace of $\mathbb{P}(\wedge^{2}V)$ consisting of bivectors of rank at most $2s$ where $2s=v-1-2q$. There are noncommutative crepant resolutions for $\operatorname{Pf}_{s}$ and $\operatorname{Pf}_{q}$ respectively denoted by $A$ and $B$, see [ŠVdB17, Section 6]. Note this gives noncommutative structure sheaves $(\operatorname{Pf}_{s},A)$ and $(\operatorname{Pf}_{q},B)$. These are homological projective duals of one another, cf. [RS19]. Let $L$ be a generic subspace of $\wedge^{2}V^{\vee}$ with dimension $sv$. Moreover, the sheaves $A,B$ restrict to give noncommutative crepant resolutions of the varieties $\operatorname{Pf}_{s}\cap\mathbb{P}(L^{\perp}),\operatorname{Pf}_{q}\cap\mathbb{P}(L)$ such that $D^{b}_{\operatorname{coh}}(A|_{\mathbb{P}(L^{\perp})})$ and $D^{b}_{\operatorname{coh}}(B|_{\mathbb{P}(L)})$ are equivalent categories. This is [RS19, Theorem 1.1]. The reader is encouraged to see [RS19] for further details; including the case of even-dimensional vector spaces.

### v3

# A note on generation and descent for derived categories of noncommutative schemes

@@ Abstract.
This work demonstrates classical generation is preserved by the derived pushforward along the canonical morphism of a noncommutative scheme to its underlying scheme. There are intriguing examples illustrating this phenomenon, particularly from noncommutative resolutions, categorical resolutions, and homological projective duality. Additionally, we establish that the Krull dimension of a variety over a field is a lower bound for the Rouquier dimension of the bounded derived category associated with a noncommutative scheme on it. This is an extension of a classical result of Rouquier to the noncommutative context.
@@
Derived categories, strong generation, Rouquier dimension, noncommutative scheme, noncommutative crepant resolutions, categorical resolution of singularities, homological projective duality, Azumaya algebra
@@ 1. Introduction
In this note, we examine the behavior of generation along the derived pushforward of the canonical morphism from a noncommutative scheme to its underlying scheme. As a result, our investigation introduces new methods coming from noncommutative homological techniques for establishing upper bounds on the Rouquier dimension within triangulated categories relevant to the commutative world.
@@
There has been recent advancements that illuminate sufficient conditions under which $D^{b}_{\operatorname{coh}}(X)$ possesses such objects: for strong generators, any Noetherian quasi-excellent separated scheme of finite Krull dimension [Aok21]; for classical generators, any Noetherian $J\textrm{-}2$ scheme [ELS20]. Albeit being aware of the existence of such objects, explicit descriptions are limited to a few instances: smooth quasi-projective schemes over a field [Rou08], Frobenius pushforwards on a compact generator for Noetherian schemes of prime characteristic [BIL+23], singular varieties that admit a resolution of singularities [Lan23, DL24], and objects arising from module categories [DLT23].
@@
Recent attention has been drawn to the case where $X$ is a projective curve [BD11, BDG17b, HH23, BD17] with significant progress made in establishing upper bounds on the Rouquier dimension of its bounded derived category. In particular, [BDG17b] identified an essentially surjective functor $D^{b}_{\operatorname{coh}}(\mathbb{X})\to D^{b}_{\operatorname{coh}}(X)$, where $\mathbb{X}$ is a noncommutative projective curve, yielding a categorical resolution of $X$. In the context of this work, a noncommutative scheme is defined as a pair $(X,\mathcal{A})$, where $X$ is a Noetherian scheme, and $\mathcal{A}$ is a coherent $\mathcal{A}$-algebra [YZ06, BDG17a]. Additional background information can be found in Section 3.1. For alternative approaches to noncommutative algebraic geometry, refer to [Orl16, KR00, Lau03, BRS+16].
In a closely related context, [ELS20] demonstrated that the bounded derived category of coherent $\mathcal{A}$-modules, denoted $D^{b}_{\operatorname{coh}}(\mathcal{A})$, admits a classical generator whenever $\mathcal{A}$ is a coherent $\mathcal{O}_{X}$-algebra over a Noetherian $J\textrm{-}2$ scheme. Our work is motivated by two key aspects: the preservation of generation along a derived pushforward of a proper surjective morphism and the geometry of a noncommutative scheme. This represents a blend of ideas from [Lan23, DL24, ELS20, BDG17a] and leads us to our first result: an investigation into the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative scheme to its underlying scheme. The following is Theorem 3.1.
@@
There are interesting examples where Theorem A can be studied, as explored in more detail in Section 4. In particular, these include noncommutative resolutions (NCR), categorical resolution of singularities, and homological projective duality. These examples not only showcase interesting applications of noncommutative schemes but also provide valuable strategies offering upper bounds for Rouquier dimension within the commutative setting. Furthermore, we introduce a construction that yields examples of NCRs, where the obtained bounds indicate that the Rouquier dimension is no larger than the global dimension of a specific endomorphism ring. This is illustrated in Lemma 4.2.3 and discussed further in Remark 4.2.4. The subsequent Corollary 3.2 encapsulates the results presented below.
@@ 2. Generation
    [BIL+23, Corollary 3.9] If $X$ is a Noetherian $F$-finite scheme, $G$ is a compact generator, and $e\gg 0$, then $F_{\ast}^{e}G$ is a classical generator for $D^{b}_{\operatorname{coh}}(X)$.
@@ 3. Derived descent
In this section, we investigate the preservation of classical generators through the derived pushforward along the canonical morphism from a noncommutative scheme to its underlying scheme. The subsequent is our main result.
@@
Before delving into the proofs of our results, we take a brief detour to explore the essential background and machinery related to noncommutative schemes.
3.1. Noncommutative schemes
@@
    A noncommutative scheme is a pair $(Y,\mathcal{B})$ where $Y$ is a scheme and $\mathcal{B}$ is a quasi-coherent sheaf of $\mathcal{O}_{Y}$-algebras.
@@
If $X$ is a scheme, then the pair $(X,\mathcal{O}_{X})$ is a noncommutative scheme. Assume $\mathcal{A}$ is a quasi-coherent $\mathcal{O}_{X}$-algebra. Consider the morphism of noncommutative schemes:
@@
Let $X$ be a variety over a field. Consider a sheaf of Azumaya algebras $\mathcal{A}$ on $X$, i.e., a locally finite free sheaf on $X$ that is étale locally a matrix algebra. This sheaf corresponds to an element in the Brauer group of $X$. For a treatment of the functorial properties of the noncommutative scheme $(X,\mathcal{A})$ and its bounded derived category, refer to [Kuz06, Appendix D]. The pair $(X,\mathcal{A})$ is called a Azumaya variety.
@@ 3.2. Results
To exhibit Theorem 3.1, we start by checking for exactness of the canonical morphism of a noncommutative scheme to its underlying scheme. The following statement might be clear to experts, but we include it for the sake of coherence.
@@
${{\mathcal{A}}}$${{j^{\ast}\mathcal{A}}}$${{\mathcal{O}_{X}}}$${{j_{\ast}\mathcal{O}_{Z}}.}$$\scriptstyle{\phi_{1}}$$\scriptstyle{\phi_{2}}$$\scriptstyle{\theta}$$\scriptstyle{\pi}$
@@ 4.4. Homological projective duality
Suppose $V$ is a $k$-vector space of odd dimension $v$. For each even number $0\leq 2q<v$, we have a Pfaffian variety $\operatorname{Pf}_{q}$ contained in $\mathbb{P}(\wedge^{2}V^{\vee})$ consisting of all $2$-forms on $V$ whose rank is at most $2q$. This variety is not a complete intersection and usually is singular. The projective dual of $\operatorname{Pf}_{q}$ is another Pfaffian variety whose locus is the subspace of $\mathbb{P}(\wedge^{2}V)$ consisting of bivectors of rank at most $2s$ where $2s=v-1-2q$. There are noncommutative crepant resolutions for $\operatorname{Pf}_{s}$ and $\operatorname{Pf}_{q}$ respectively denoted by $A$ and $B$, see [ŠVdB17, Section 6]. Note this gives noncommutative schemes $(\operatorname{Pf}_{s},A)$ and $(\operatorname{Pf}_{q},B)$. These are homological projective duals of one another, cf. [RS19]. Let $L$ be a generic subspace of $\wedge^{2}V^{\vee}$ with dimension $sv$. Moreover, the sheaves $A,B$ restrict to give noncommutative crepant resolutions of the varieties $\operatorname{Pf}_{s}\cap\mathbb{P}(L^{\perp}),\operatorname{Pf}_{q}\cap\mathbb{P}(L)$ such that $D^{b}_{\operatorname{coh}}(A|_{\mathbb{P}(L^{\perp})})$ and $D^{b}_{\operatorname{coh}}(B|_{\mathbb{P}(L)})$ are equivalent categories. This is [RS19, Theorem 1.1]. The reader is encouraged to see [RS19] for further details; including the case of even-dimensional vector spaces.